OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -11, the last term is -171,875, and the common ratio is -5?

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

For a finite geometric sum, recall the formula \(\displaystyle \sum_{k=1}^{n} ar^{k-1} = a\dfrac{1-r^n}{1-r}\), where \(a\) is the first term, \(n\) is the number of terms in our series, and \(r\) is the common ratio. In this case, you're given that a=-11, n=7, and r=-5. What do you get when you plug everything into that formula? Can you take things from here? :-)

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

-143,231 -36,047 144,177 716,144

OpenStudy (anonymous):

@ChristopherToni

OpenStudy (anonymous):

What do you get when you evaluate \(\displaystyle -11\cdot\frac{1-(-5)^7}{1-(-5)}\) with your calculator?

OpenStudy (anonymous):

my calculator says error

OpenStudy (anonymous):

-859441

OpenStudy (anonymous):

@ChristopherToni